Week 31
This presentation is the property of its rightful owner.
Sponsored Links
1 / 35

Week 31 PowerPoint PPT Presentation


  • 64 Views
  • Uploaded on
  • Presentation posted in: General

Week 31. Summary week 31 02.08.-09.08. G . Arduini & O . Brüning. Week 31. Operation with Multi-bunch Injection since Thursday 29.07: 13 fills with up-to 120 nb -1 per fill  109 h of luminosity production

Download Presentation

Week 31

An Image/Link below is provided (as is) to download presentation

Download Policy: Content on the Website is provided to you AS IS for your information and personal use and may not be sold / licensed / shared on other websites without getting consent from its author.While downloading, if for some reason you are not able to download a presentation, the publisher may have deleted the file from their server.


- - - - - - - - - - - - - - - - - - - - - - - - - - E N D - - - - - - - - - - - - - - - - - - - - - - - - - -

Presentation Transcript


Week 31

Week 31

Summary week 31

02.08.-09.08.

G. Arduini & O. Brüning

8:30 meeting


Week 311

Week 31

  • Operation with Multi-bunch Injection since Thursday 29.07:

    • 13 fills with up-to 120 nb-1 per fill  109 h of luminosity production

       ca. one fill per day; peak luminosity: L = 4 1030 cm-2 sec-1

    • average fill time of ca. 9 hours (10, 9, 14, 1, 15, 7, 2, 17, 15, 0.5, 2, 14, 2.5)

    • integrated luminosity: 730 nb-1 (40, 60, 80, 8, 100, 60, 20, 120, 90, 8, 23, 95,25)

    •  ca. 56 nb-1 /fill;  average luminosity per fill: L = 1.7 1030 cm-2 sec-1

       ca. 40% efficiency (wrt to scheduled time);  total luminosity ca. 1080 nb-1

8:30 meeting


Week 312

Week 31

8:30 meeting


Week 313

Week 31

For periods without studies: ca. 63% efficiency relative to available beam time

ca. 45% efficiency per scheduled time

8:30 meeting


Monday 2 8

Monday 2.8.

  • Analysis of the noisy spectra observed at injection and flat-top:

    • The peaks observed are exact harmonics of 50 Hz extending up to 4 kHz

    • Instrumental?

    • The noise seems to be lower in the BBQ pick-up for on-demand measurements

8:30 meeting


Tuesday 3 8

Tuesday 3.8.

  • Test Ramp for operation optimization with 2 nominal bunches:

    • Beams lost during squeeze at b* = 7m

  • Goal for test ramp with 2 bunches for operation optimizations:

    • A: Damper and tune feedback optimization:

      • Test of the Q feedback with different PC regulation loop to reduce noise ✔

      • Transverse feedback tests at injection (0.5 hours) Not yet done

      • Hump investigations (~1.5 hours) Not yet done

    • B: Ramp and squeeze optimization:

      • Incorporate Q and Orbit FB feed forward into ramp. ✔

      • Establish better end-squeeze to collision sequence/settings. Not yet done

      • Clean orbit in stable beams (ALICE H plane,..). Not yet done

      • Check TCTs. Not yet done

      • Collimator loss maps. Not yet done

      • Beam dump validation (Dump with RF OFF) Not yet done

8:30 meeting


Wednesday 4 8

Wednesday 4.8.

  • Tuesday to Wednesday night:

    • Physics fill with 25 nominal bunches with inverted sign of tune split;

    • Problems related to losses during over-injection into LHC

    • Initial peak luminosity of 2.8 1030 cm-2 sec-1; fill lost due to patrol box in IR1

  • Wednesday morning: optimization of losses during injection:

    • Single bunch has very good trajectory and no losses in the TL.

    • Over-injection with 4 bunches leads to losses that dump the beam  trajectory changes by ca. 0.2mm over all cases (1 to 4 bunch injection); trimmed SPSRF frequency (10 Hz);

    • Decided to correct trajectory in both lines back to reference using

      upstream correctors (0.6mm rms) (origin of change not understood?).

    • Revalidation of TCDI settings.

    • Change of some BLM thresholds;

8:30 meeting


Tuesday 3 81

Tuesday 3.8.

  • Beam intensities over fill– Losses for Beam2 higher than for Beam1 after inverted sign of tune split

8:30 meeting


Wednesday 4 81

Wednesday 4.8.

  • Afternoon: Preparation of new physics fill:

    • Physics fill with 25 nominal bunches; No tune split between Beam1 and Beam2

    • 15:45: Starting ramp: e1nx = 2 mm; e1ny = 3 mm; e2nx = 2.7 mm; e2ny = 5.3 mm;

    • 17:00: Starting Squeeze;

    • 17:40: Stable Beams; Initial peak luminosity of 3.0 1030 cm-2 sec-1; Beam lifetimes of 5h going above 25 hours after few minutes and rising to 125h after a couple of hours.

    • 19:30: Roman Pots at 25 sigma;

8:30 meeting


Wednesday thursday 4 5 8

Wednesday-Thursday 4-5.8.

  • End of physics at ~11:00 after 18.5 hours

~120 nb-1

8:30 meeting


Thursday 5 8

Thursday 5.8.

  • Go to adjust mode for end of coast study

  • Swap H/V tunes for B2 (crossing coupling resonance) in order to investigate the feasibility of working point with similar to RHIC in collision. Damper can stand the tune swap but not completely transparent (dip in lifetime and tiny losses)

8:30 meeting


Thursday 5 81

Thursday 5.8.

  • Afternoon: Second attempt for a test fill for ramp and squeeze optimization

    • (2 on 2 bunches)

  • 20:00 RQ10.B2 power converter trip (water fault again)  Beam dump

    • Incorporate Q and Orbit FB feed forward into ramp. ✔

    • Establish better end-squeeze to collision sequence/settings.✔

    • Clean orbit in stable beams (ALICE H plane,..). ✔

    • Check TCTs. Not yet done

    • Collimator loss maps. Not yet done

    • Beam dump validation (Dump with RF OFF) Not yet done

8:30 meeting


Thursday 5 82

Thursday 5.8.

  • New beam process for collision (Mike):

  • All went smoothly ! Non-closure of the knobs 30 microns in V, 60 microns in H. OK !  To be implemented in the sequence

8:30 meeting


Thursday 5 83

Thursday 5.8.

  • Access request from CMS to fix serious problem on their detector

  • Profited from the access for CMS to fix:

    • problem on MKB (vacuum interlock configuration)

    • Module replacement for fixing the spurious interlocks on Q10.R8

  • Accesses finished by 22:45

  • Tedious injection process because of an intermittent problem with an 80 MHz cavity in the PS.

8:30 meeting


Friday 6 8

Friday 6.8.

  • 04:00: Fill 1263

    • Physics fill with 25 nominal bunches; No tune split and DQ = +2 10-3;

    • Emittances: ex1 = 3.4 mm; ey1 = 3.0mm; ex2 = 3.0 mm; ey2 = 3.2mm;

    • Initial peak luminosity of 2.7 1030 cm-2 sec-1 (ca. 10% less then previous fill);

    • Beam lifetimes initially 5h and above 25 hours after a couple of hours.

    • Terminated at 19:15 Generated integrated luminosity of ca. 90 nb-1

    • total luminosity delivered ca. 980 nb-1

8:30 meeting


Friday 6 81

Friday 6.8.

  • Feedback

    during

    ramp and

    squeeze:

8:30 meeting


Friday 6 82

Friday 6.8.

  • Relative

    bunch

    intensity

    drops

    for

    Beam1

    and

    Beam2:

    Fill 1263

GiuliaPapotti (BE/OP/LHC)

8:30 meeting


Friday 6 83

Friday 6.8.

  • Relative

    bunch

    intensity

    for

    Beam1

    and

    Beam2:

    Fill 1263

    no Q split &

    DQ = 2 10-3

    versus

    Fill 1262

    noQ split

giulia papotti (BE/OP/LHC)

8:30 meeting


Friday 6 84

Friday 6.8.

Relative

bunch

intensity

drops

for

Beam1

and

Beam2:

Fill 1263

no Q split &

DQ = 2 10-3

versus

Fill 1257

with Q split

giulia papotti (BE/OP/LHC)

8:30 meeting


Friday 6 85

Friday 6.8.

  • Luminosity over 16 hours:

     ca. 50% drop over 15 hours

~90 nb-1

8:30 meeting


Friday 6 86

Friday 6.8.

  •  Champagne!!!

  • Fill terminated at 19:00:  ca. 1000 nb-1 ± 10%

8:30 meeting


Friday 6 87

Friday 6.8.

  • 21:00: Preparing new fill for the night: Fill 1264

    • 21:00: injecting pilot and setup (some delays: mastership, phasing, SPS beam quality [emittance and intensity])  start ramp around midnight

    • Emittances: ex1 = 2.4 mm; ey1 = 2.3mm; ex2 = 2.4 mm; ey2 = 3.6mm;

    • 01:42: Stable beams:

    • Initial peak luminosity of 3.1 1030 cm-2 sec-1;

    • Beam lifetimes initially 5h and above 25 hours after a couple of hours.

  • 02:15: Beam dump (BLM in Pt4)

    • Start pre-cycle  machine ready for new injection at

8:30 meeting


Saturday 7 8

Saturday 7.8.

  • Initial Luminosities

8:30 meeting


Saturday 7 81

Saturday 7.8.

  • Lifetimes rising from 5h to above 25 hours over one hour:

8:30 meeting


Saturday 7 82

Saturday 7.8.

  • Fill #1264 dump at 02:14:38: BLM @Q11.L4B1 (2.5 ms threshold triggered) .

  • The comparison with a BLM at TCP.L7 indicates that the loss is real, similar time behaviour


Saturday 7 83

Saturday 7.8.

  • 04:15: Problems resetting the machine

    • Can't close the switch of RQTF.A67B2 called piquet and preparing for access in UA67 (but could closed in the end by M. Zerlauth by doing "Reset System" )

  • 05:24: Sector 67 tripped again:

    • Quench loop of Q8.R6 which opened again.

    • As before, a disturbance in the quench signal opens the quench loop.

    • Diagnosis later points to the power converter as origin for the fault

  • 06:15: A local perturbation in Pt2 stopped the filters of the compensator; TI advises to stop sectors S12 and S23 + the ALICE magnets

  •  Spent all day in access for QPS and FGC interventions!

8:30 meeting


Sunday 7 8

Sunday 7.8.

  • 20:00: Preparing fill for the night: Fill 1266

    • Emittances: ex1 = 2.2mm; ey1 = 2.3mm; ex2 = 2.0mm; ey2 = 2.5mm;

    • 23:00: Stable beams:

    • Initial peak luminosity of 4 1030 cm-2 sec-1;

    • Beam lifetimes initially 5h and above 25 hours after a couple of hours.

    • Beam dumped after 2h! This time by BLM in IR5.

  • 01:00: Preparing new fill for night:

    • Emittances: ex1 = 2.5mm; ey1 = 2.5mm; ex2 = 2.6mm; ey2 = 4.0mm;

    • 05:30: Stable beams:

    • Initial peak luminosity of 3 1030 cm-2 sec-1;

    • Beam lifetimes initially 5h and above 25 hours after a couple of hours.

    • 18:50: Loss of Cryo in S34 (loss of communication)  beam dump after 14 hours.

8:30 meeting


07 08 2010

07/08/2010

  • Relative losses for fill 1264 and 1266.

    • Both with NO tune split.

    • Both dumped after one hour due to BLM reading

  • 1266 initial losses lower than in any other fill

G. Papotti


Sunday 8 8

Sunday 8.8.

  • 21:00: Short Hump study at 450 GeV

    • No clear evidence yet on effect of offset in cavities on Hump activity.

  • 23:00: Preparing fill for the night:

    • Emittances: ex1 = 2.7mm; ey1 = 2.9mm; ex2 = 2.8mm; ey2 = 3.9mm;

    • 01:30: Stable beams:

    • Initial peak luminosity of 3 1030 cm-2 sec-1;

    • Beam lifetimes initially 5h and above 25 hours after a couple of hours.

    • 04:00: Beams dumped: quench loop of the RB.A67

8:30 meeting


Saturday 7 84

Saturday 7.8.

  • Plan for the coming days:

    • Luminosity fills

    • Next week: Test ramp with 2 bunches:

      • Check TCTs.

      • Collimator loss maps.

      • Beam dump validation (Dump with RF OFF)

      • Implementation of new beam process for establishing collisions.

8:30 meeting


Pending

Pending

  • Tests:

    • New BLM Firmware tests

    • Delay of the PM trigger of the BLM system to be set back  BLM team ready only next week

    • Loss maps, orbit clean-up and TCT verifications

  • Requiring access:

    • Repair COD MCBH.18R7.B2

    • CV need to replace a motorization of the water valve of the sector 56 - not urgent

      • Time estimate ~3 hours

    • Access team to replace oxydized connectors for PAD-PX24 in case of stop (3 hours)

    • Magnetic probes in the machine for the studies related to the hump - access in point 5 (GA, LW) – 3 hours


Monday 9 8

Monday 9.8.

  • Open Issues:

    • Clogging of CMS cryo filters not yet understood

    • QPS boards: several boards had to switch from channel A to B.

    • BLM racks (in buildings SR1 and SR5 and both at position BY02): would like to correct the way the power cable for the multi-plug arrives inside the rack.

    • Ion pumps in dump line MKB (B2 only 4 out of 8 pumps working; few days intervention, next TS)

    • Origin – cause of the beam losses that terminated 2 of the best runs (beam-beam?; coherent versus incoherent?).

    • Bunch-to-Bunch diagnostics would be very helpful!

  • Week 32:

    • Malika and Jörg

8:30 meeting


Monday 2 81

Monday 2.8.

  • 12:35: Lost access chain in point 2 (PX24 PAD)  Dump after ~14.5 hours

~100 nb-1

8:30 meeting


Tuesday 3 82

Tuesday 3.8.

  • Sector 23 Quadrupoles tripped at 07:39 after 7 hours and 20 mins in coast.

  • Cause being investigated

  • CMS preparing to ramp down the solenoid for cryo intervention

~50 nb-1


Tuesday 3 83

Tuesday 3.8.

  • Beam2 losses during over injection:

     Losses for Beam2 too high for operation

8:30 meeting


  • Login